OpenStudy (arivillafana):

I WILL FAN,MEDAL,& TESTIMONIAL -Find e, if a=8, d=22, b=4 -Find d, if b=6, e=15, a=8 -Find a, if f=32, c=12, d=24 -Find c, if f=15, b=7, e=3 -Find f, if c=2 1/2, d=24, a=10 -Find b, if c=5, e=7, f=9 -Find a, if b=11, e=8, d=6

OpenStudy (anonymous):

i bet we are missing something

OpenStudy (anonymous):

an expression involving e i hope,

OpenStudy (arivillafana):

@satellite73 there are triangles, but i cant send a picture its a weird file, i can send you the actual thing ?

OpenStudy (anonymous):

by all means post something

OpenStudy (arivillafana):

@satellite73 there

OpenStudy (anonymous):

oh wow ok it is mostly solving ratios

OpenStudy (arivillafana):

yeah pretty much, all I have solved is the two riddle up top

OpenStudy (anonymous):

which ones have you done? i don't want to repeat any

OpenStudy (arivillafana):

ive only done the riddles up top

OpenStudy (anonymous):

ok lets do the first one

OpenStudy (anonymous):

damn i labelled them all wrong

OpenStudy (arivillafana):

yeah I was confused lol

OpenStudy (anonymous):

\[\frac{e}{22}=\frac{4}{8}\] that's better

OpenStudy (anonymous):

which really just means \(\frac{e}{22}=\frac{1}{2}\) i.e. \(e\) is half of \(22\)

OpenStudy (arivillafana):

i feel so dumb right now

OpenStudy (anonymous):

did you understand what i wrote?

OpenStudy (arivillafana):

not really

OpenStudy (anonymous):

we can go slower if you like

OpenStudy (arivillafana):

yes, that'd be helpful ((:

OpenStudy (anonymous):

ok we have two triangles that are similar that means a) all the angles are the same and b) the ratios of the corresponding sides are the same you can forget about the angles for this one just know that the ratios of the sides are the same let me draw the picture correctly this time

OpenStudy (arivillafana):

okay

OpenStudy (anonymous):



OpenStudy (anonymous):

so for example \[\frac{a}{b}=\frac{d}{e}\] do you see how i got that ?

OpenStudy (arivillafana):

i think so

OpenStudy (anonymous):

right side over the bottom on the left, right sides over the bottom on the right

OpenStudy (anonymous):

lets write another one \[\frac{b}{c}=\frac{e}{f}\]

OpenStudy (arivillafana):

okay I get how you got that

OpenStudy (anonymous):

bottom over left on the right, bottom over left on the left get the idea?

OpenStudy (arivillafana):

yes

OpenStudy (anonymous):

there are more we can write

OpenStudy (arivillafana):

okay

OpenStudy (anonymous):

\[\frac{c}{a}=\frac{f}{d}\] for example

OpenStudy (arivillafana):

so for the first one, would e=4 ?

OpenStudy (anonymous):

and so on and so one because we can flip any of these \[\frac{c}{a}=\frac{f}{d}\iff \frac{a}{c}=\frac{d}{f}\]

OpenStudy (anonymous):

lets make a ratio that includes \(a, d, b, e\) because we know three out of those four numbers

OpenStudy (arivillafana):

okay

OpenStudy (anonymous):

\[\frac{b}{a}=\frac{e}{d}\] do you see that one?

OpenStudy (arivillafana):

yes

OpenStudy (anonymous):

now lets plug in the numbers we know for \(a,b,d\)

OpenStudy (anonymous):

\[\frac{4}{8}=\frac{e}{22}\] uses

OpenStudy (anonymous):

you see what i did? plugged in the numbers you were told for \(a=8,b=4,d=22\)

OpenStudy (arivillafana):

yes

OpenStudy (anonymous):

final job is to solve this for \(e\) do you know how to do that ? (no is a fine answer, i am just asking)

OpenStudy (arivillafana):

well, since b is similar to e, wouldn't e= 4 also ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if they were "congruent' it would be 4

OpenStudy (arivillafana):

oh

OpenStudy (anonymous):

you gotta solve this \[\large \frac{4}{8}=\frac{e}{22}\] for \(e\)

OpenStudy (anonymous):

if the sides of both triangles were the same, you would have nothing to do in these problems you have to solve the ratios if you do not know how i will tell you

OpenStudy (arivillafana):

how would i solve the ratio to find e ?

OpenStudy (anonymous):

is it clear that 4 is half of 8?

OpenStudy (arivillafana):

yes

OpenStudy (anonymous):

that means \(e\) must be half of 22

OpenStudy (arivillafana):

so e= 11 ?

OpenStudy (anonymous):

right since \[\frac{4}{8}=\frac{11}{22}=\frac{1}{2}\]

OpenStudy (anonymous):

wanna do another?

OpenStudy (arivillafana):

yes, can we do: Find f, if c=2 1/2, d=24, a=10

OpenStudy (anonymous):

ok we need a ratio involving \(f, c, d, a\)

OpenStudy (arivillafana):

okayy.. so would it be 2 1/2 f ---- --- 24 10

OpenStudy (anonymous):

hold on let me check

OpenStudy (arivillafana):

okay

OpenStudy (anonymous):

no i don't think so let me write one

OpenStudy (arivillafana):

okay

OpenStudy (anonymous):

or let me ask you this' what is \(\frac{f}{d}\) equal to in the other triangle (just with the letters, forget the numbers)

OpenStudy (arivillafana):

c/a ?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

now lets put the numbers in

OpenStudy (arivillafana):

sorry i dont know how you make into a fraction lol so i had to do that

OpenStudy (anonymous):

i will do it, don't fret

OpenStudy (anonymous):

\[\frac{f}{24}=\frac{2\tfrac{1}{2}}{10}\]

OpenStudy (anonymous):

i just plugged in the numbers, hope i did it correctly it is hard to keep track

OpenStudy (arivillafana):

ohhh okay i see what i did wrong the first time

OpenStudy (anonymous):

now we have to solve this for \(f\) right?

OpenStudy (arivillafana):

yes

OpenStudy (anonymous):

i.e. we have to get rid of the \(24\) in the denominator by multiplying both sides by \(24\)

OpenStudy (anonymous):

\[\frac{f}{24}=\frac{2\tfrac{1}{2}}{10}\\ f=24\times \frac{2\tfrac{1}{2}}{10}\]

OpenStudy (anonymous):

mixed numbers suck to work with unless you can do \(24\times 2\tfrac{1}{2}\) in your head

OpenStudy (arivillafana):

I cant lol

OpenStudy (anonymous):

actually you can two times twenty four is forty eight, half of twenty four is twelve and forty eight plus twelve is sixty but lets do it the math teacher way

OpenStudy (arivillafana):

haha okay

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

turn \(2\tfrac{1}{2}\) in to an improper fraction \[2\tfrac{1}{2}=\frac{5}{2}\] then \[24\times \frac{5}{2}=12\times 5=60\] as promised